[Brahms-l] Shift report 20050316 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Mar 16 2005 - 00:01:53 EST
Shifter leader was kh
Shifters were 


The only beam we had during this shift was from about 16:15 to 18:00 where we had injection energy beam that was left for data taking during that time.

At that point they began to try for 62GeV beam again.  There was a Quench Link Interlock; then various aborted stores;  then communications problem with 5 o'clock frontends.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Mar 16 00:02:24 2005

This archive was generated by hypermail 2.1.8 : Wed Mar 16 2005 - 00:02:37 EST